7: If ?(?) = 0.35 , ?(?) = 0.25, and the events ? and ? are disjoint (mutually exclusive), then ?(? ?? ?) = _________.
Answer #1

Solution:

P(A) = 0.35

P(B) = 0.25

P(A or B) = P(A) + P(B)

= 0.35 + 0.25 = 0.6

P(A or B) = 0.6